Need the rule (slope intercept equation) for the following points (1,1), (3,2), (5,5), (7,4)
Also how do you determine the rule given the above points?

Answer by ewatrrr(24785)

 
Hi,
(7,4)
(5,5) m = -1/2 y = (-1/2)x + 15/2
(5,5)
(3,2) m = 3/2
(3,2)
(1,1) m = 1/2 y = (1/2)x + 1/2
These points are not ALL on the same line, if they were..the behavior of
change (change in y/change in x) going from point to point would be the same